Describe the sampling distribution of p. Assume the size of the population is 30,000. n=800​, p=0.4 Choose the phrase that best describes the shape of the sampling distribution of p below.

A.

Not normal because

n≤0.05N

and np(1−p)≥10.

B.Not normal because

n≤0.05N

and np(1−p)<10.

C.Approximately normal because

n≤0.05N

and np(1−p)<10.

D.Approximately normal because

n≤0.05N

and np(1−p)≥10.

Determine the standard deviation of the sampling distribution of p.
Answer #1

Solution

Given that,

p = 0.4

1 - p = 1 - 0.4 = 0.6

n = 800

D.Approximately normal because n≤0.05N and np(1−p)≥10.

\mu\hat p = p = 0.4

\sigma\hat p  \sqrt[p ( 1 - p ) / n] = \sqrt [(0.4 * 0.6) / 800 ] = 0.0173